OpenStudy (anonymous):

I know its simple but I keep getting different answers :( What is the simplified form of 2/x^2-x - 1/x ? Really appreciate it if I could get some help :)

OpenStudy (dumbcow):

\[\frac{2}{x^{2}}-x-\frac{1}{x}\] is that it?

OpenStudy (anonymous):

no the (-x) is the denominator with x^2. its x^2-x :)

OpenStudy (dumbcow):

oh ok...in that case i would factor out an x \[\frac{2}{x^{2}-x}-\frac{1}{x} = \frac{2}{x(x-1)}-\frac{1}{x} = \frac{2}{x(x-1)}-\frac{1(x-1)}{x(x-1)}=\frac{3-x}{x(x-1)}\]

OpenStudy (anonymous):

Thanks so much! :D

OpenStudy (dumbcow):

your welcome